OpenStudy (anonymous):

Ben is 3 years older than Will and Jen is two times older than Ben the sum of their ages is 41 how would I set this up?

OpenStudy (kropot72):

Let the age of Ben be B, the age of Will be W and the age of Jen be J. W = B - 3 J = 2B B + J + W = 41 ...........(1) Plugging the above values in terms of B for J and W into equation (1) we get: B + 2B + B - 3 = 41 ...........(2) Now just solve equation (2) to find the value of B.

OpenStudy (anonymous):

Thank you kropot72

OpenStudy (kropot72):

You're welcome :)